OpenStudy (anonymous):

use the method of substitution to solve 2x-y=8 9x-4y=37

OpenStudy (anonymous):

first question: do you know what the method of substitution is. if not please look at the Khan academy videos on the subject.

OpenStudy (anonymous):

2x - y = 8 ==> -y = 8 - 2x ==> y = 2x - 8 now you sub 2x - 8 in for y in the 2nd equation... 9x - 4y = 37 9x - 4(2x - 8) = 37 You then distribute the -4 through the parenthesis and solve for x Once you have x solved you can substitute that value into either one of the original equations and solve for y. You can then check your answer by subbing in your known variables into either one of the original equations and if the solution comes out equal to each other, then you have done it correctly :)

OpenStudy (anonymous):

ok so far this is what i got 9x-4(2x-8)=37 9x-8x-32=37

OpenStudy (anonymous):

2x-y=8 9x-4y=37 ---------- (-4)(2x-y) = (-4)(8) 9x-4y = 37 ----------------- -8x+4y = -32 9x-4y = 37 ------------- // add vertically (-8+9)x+(4-4)y = (-32+37) x +0y = 5 x = 5 plug into either equation to get the y value.

OpenStudy (anonymous):

jasmineplayhouse check what you just did above again...And remember that a negative times a negative equals a positive.

OpenStudy (anonymous):

ok i sorry but i am still confused

OpenStudy (anonymous):

Let me show you... 9x - 4(2x - 8) = 37 (distribute the -4 through the parenthesis) 9x - 8x + 32 = 37 (combine like terms) x + 32 = 37 (subtract 32 from both sides to get x by itself) x + 32 - 32 = 37 - 32 x = 5 now sub 5 in for x in either of the original equations... 2x - y = 8 2(5) - y = 8 10 - y = 8 10 - 8 = y 2 = y now check... 9x - 4y = 37 9(5) - 4(2) = 37 45 - 8 = 37 37 = 37 (correct) Do you understand everything ? If not, ask questions and I will try to answer them.

OpenStudy (anonymous):

ok i substitute x with 5 and y with 2

OpenStudy (anonymous):

yes...because we worked it out and x = 5 and y = 2. It was checked and found to be correct.

OpenStudy (anonymous):

so the order pair is 37=37

OpenStudy (anonymous):

ok (2,5) would be the order pair

OpenStudy (anonymous):

No...the ordered pair is (x,y) = (5,2). The 37's were the result of me checking my work. If the end result was not equal, for instance if it was 37 = 35, then I know the answer would be wrong and I would have made a mistake.However, they equaled so it was correct.

OpenStudy (anonymous):

ok i got it! thanks

OpenStudy (anonymous):

your welcome...thats what I am here for.